OpenStudy (anonymous):

What is the equation of a line that passes through the point (2, 5) and is perpendicular to a line with a slope of -2

OpenStudy (anonymous):

ok i get the slope is the reciprocal of -2 so its 2 y=2x +b how do i find b......it says it can be +4 or +5

OpenStudy (triciaal):

use the same process as before perpendicular lines you can find the slope you have 2 points (x, y) and (2, 5)

OpenStudy (aum):

negative reciprocal of -2 is 1/2

OpenStudy (anonymous):

5=1/2(2)+b like this

OpenStudy (aum):

yes.

OpenStudy (anonymous):

so its 4

OpenStudy (aum):

b = 4 y = 1/2x + 4

OpenStudy (anonymous):

y=1/2x+4

OpenStudy (anonymous):

Awwww bless you guys.

OpenStudy (triciaal):

ok i get the slope is the reciprocal of -2 so its 2 no the slope is -1/(-2/1) so it is -1*(-1/2) to get 1/2

OpenStudy (aum):

yw.

OpenStudy (anonymous):

What is the equation of a line that passes through the point (0, 2) and is perpendicular to a line with a slope of 3

OpenStudy (anonymous):

so i got y=-1/3x+2